# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ Abstract
We study the problem of gradient descent learning of a single-index target function $f_{*}(\boldsymbol{x})=\textstyle\sigma_{*}\mathopen{}\mathclose{{}\left(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle}\right)$ under isotropic Gaussian data in $\mathbb{R}^{d}$, where the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ is an unknown degree $q$ polynomial with information exponent $p$ (defined as the lowest degree in the Hermite expansion). Prior works showed that gradient-based training of neural networks can learn this target with $n\gtrsim d^{\Theta(p)}$ samples, and such statistical complexity is predicted to be necessary by the correlational statistical query lower bound. Surprisingly, we prove that a two-layer neural network optimized by an SGD-based algorithm learns $f_{*}$ of arbitrary polynomial link function with a sample and runtime complexity of $n\asymp T\asymp C(q)\cdot d\mathrm{polylog}d$, where constant $C(q)$ only depends on the degree of $\sigma_{*}$, regardless of information exponent; this dimension dependence matches the information theoretic limit up to polylogarithmic factors. Core to our analysis is the reuse of minibatch in the gradient computation, which gives rise to higher-order information beyond correlational queries.
@@ 1 Introduction
Single-index models are a classical class of functions that capture low-dimensional structure in the learning problem. To efficiently estimate such functions, the learning algorithm should extract the relevant (one-dimensional) subspace from high-dimensional observations; hence this problem setting has been extensively studied in deep learning theory [BL20, BES+22, BBSS22, MHPG+23, WWF24], to examine the adaptivity to low-dimensional targets and benefit of representation learning in neural networks (NNs) optimized by gradient descent (GD). In this work we study the learning of a single-index target function with polynomial link function under isotropic Gaussian data:
@@
where $\varsigma_{i}$ is i.i.d. label noise, $\boldsymbol{\theta}\in\mathbb{R}^{d}$ is the direction of index features, and we assume the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ is a degree-$q$ polynomial with information exponent $p$ defined as the index of the first non-zero coefficient in the Hermite expansion (see Definition 1).
Equation (1.1) requires the estimation of the one-dimensional link function $\sigma_{*}$ and the relevant direction $\boldsymbol{\theta}$; it is known that learning is information theoretically possible with $n\gtrsim d$ training examples [Bac17, DPVLB24]. Indeed, when $\sigma_{*}$ is polynomial, such statistical complexity can be achieved up to logarithmic factors by a tailored algorithm that exploit the specific structure of the low-dimensional target [CM20]. On the other hand, for gradient-based training of two-layer NNs, existing works established a sample complexity of $n\gtrsim d^{\Theta(p)}$ [BAGJ21, BBSS22, DNGL23], which presents a gap between the information theoretic limit and what is computationally achievable by (S)GD. Such a gap is also predicted by the correlational statistical query (SQ) lower bound [DLS22, AAM23], which roughly states that for a CSQ algorithm to learn (isotropic) Gaussian single-index models using less than exponential compute, a sample size of $n\gtrsim d^{p/2}$ is necessary.
Although CSQ lower bounds are frequently cited to imply a fundamental barrier of learning via SGD (with the squared loss), strictly speaking, the CSQ model does not include empirical risk minimization with gradient descent, due to the non-adversarial noise and existence of non-correlational terms in the gradient computation. Very recently, [DTA+24] exploited higher-order terms in the gradient update arising from the reuse of the same training data, and showed that for certain link functions with high information exponent ($p>2$), two-layer NNs may still achieve weak recovery (i.e., nontrivial overlap with $\boldsymbol{\theta}$) after two GD steps with $O(d)$ batch size. While this presents evidence that GD-trained NNs can learn $f_{*}$ beyond the sample complexity suggested by the CSQ lower bound, the weak recovery statement in [DTA+24] may not translate to statistical guarantees; moreover, the class of functions where SGD can achieve vanishing generalization error is not fully characterized, as only a few specific examples of link functions are discussed.
@@
Figure 1: We train a two-layer ReLU NN (3.1) with $N=1024$ neurons using SGD with step size $\eta=1/d$ to learn a single-index target $f_{*}(\boldsymbol{x})=\mathsf{He}_{3}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)$; heatmaps are values averaged over 10 runs. $(a)$ online SGD with batch size $B=8$; $(b)$ GD on the same batch of size $n$ for $T=2^{14}$ steps. We only report weak recovery (i.e., overlap between parameters $\boldsymbol{w}$ and target $\boldsymbol{\theta}$, averaged across neurons) for online SGD since the generalization error does not drop.
Given the existence of (non-NN) algorithms that learn any single-index polynomials with $n=\tilde{O}(d)$ samples [CM20, DPVLB24], it is natural to ask if gradient-based training of NNs can achieve similar statistical efficiency for the same function class. Motivated by observations in [DTA+24] that SGD with reused batch may break the “curse of information exponent”, we aim to address the following question:
Can a two-layer NN optimized by SGD with reused batch learn arbitrary polynomial single-index models near the information-theoretic limit $n\asymp d$, regardless of the information exponent?
Empirically, the separation between one-pass (online) and multi-pass SGD is clearly observed in Figure 1, where we trained the same two-layer ReLU neural network to learn a single-index polynomial with information exponent $p=3$. We see that SGD with reused batch (Figure 1(b)) reaches low generalization error using roughly $n\asymp d$ samples, whereas online SGD fails to achieve even weak recovery with much larger sample size $n=\Omega(d^{2})$. Our main contribution is to establish this near-optimal sample complexity for two-layer NNs trained by a variant of SGD with reused training data.
@@ 1.1 Our Contributions
We answer the above question in the affirmative by showing that for (1.1) with arbitrary polynomial link function, SGD training on a natural class of shallow NNs can achieve small generalization error using polynomial compute and $n=\tilde{O}(d)$ training examples, if we employ a layer-wise optimization procedure (analogous to that in [BES+22, DLS22, AAM23]) and reuse of the same minibatch:
@@
A shallow NN with $N=\tilde{O}_{d}(1)$ neurons can learn arbitrary single-index polynomials up to small population loss: $\mathbb{E}_{\boldsymbol{x}}[(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]=o_{d,\mathbb{P}}(1)$, using $n=\tilde{O}_{d}(d)$ samples, and an SGD-based algorithm (with reused training data) minimizing the squared loss objective in $T=\tilde{O}_{d}(d)$ gradient steps.
We make the following remarks on our main result.
  • The theorem suggests that NN + SGD with reused batch can match the statistical efficiency of SQ algorithms tailored for low-dimensional polynomial regression [CM20]. Furthermore, the sample complexity is information theoretically optimal up to polylogarithmic factors, and surpasses the CSQ lower bound for $p>2$ (see Figure 2); this disproves a conjecture in [AAM23] stating that $n\asymp d^{p/2}$ is the optimal sample complexity for empirical risk minimization with SGD.
  • A key observation in our analysis is that with suitable activation function, SGD with reused batch can go beyond correlational queries and implement (a subclass of) SQ algorithms. This enables polynomial transformations to the labels that reduce the information exponent to (at most) 2, and hence optimization can escape the high-entropy “equator” at initialization in polylogarithmic time.
Upon completion of this work, we became aware of the preprint [ADK+24] showing weak recovery with similar sample complexity, also by exploiting the reuse of training data. Our work was conducted independently and simultaneously, and their contribution is not reflected in the present manuscript (beyond this paragraph).
@@
Figure 2: Sample complexity of learning single-index model where the link function $\sigma_{*}$ has degree $q$ and information exponent $p$. For the CSQ lower bound, we translate the tolerance to sample complexity using the i.i.d. concentration heuristic $\tau\approx n^{-1/2}$. We restrict ourselves to algorithms using polynomial compute; this excludes the sphere-covering procedure in [DPVLB24] or exponential-width neural network in [Bac17, TS24].
@@ 2.1 Complexity of Learning Single-index Models
We aim to learn a single-index model (1.1) where the link function $\sigma_{*}$ is a degree-$q$ polynomial with information exponent $p$ defined as follows.
Definition 1 (Information exponent [BAGJ21]).
Let $\{\mathsf{He}_{j}\}_{j=0}^{\infty}$ denote the normalized Hermite polynomials. The information exponent of $g\in L^{2}(\gamma)$, denoted by $\mathrm{IE}(g):=p\in\mathbb{N}_{+}$, is the index of the first non-zero Hermite coefficient of $g$, that is, given the Hermite expansion $g(z)=\sum_{i=0}^{\infty}\alpha_{i}\mathsf{He}_{i}(z)$, $p:=\min\{i>0:\alpha_{i}\neq 0\}$.
By definition we always have $p\leq q$. Note that $f_{*}$ contains $\Theta(d)$ parameters to be estimated, and hence information theoretically $n\gtrsim d$ samples are both sufficient and necessary for learning [MM18, BKM+19, DPVLB24]; however the sample complexity achieved by different (polynomial time) learning algorithms depends on structure of the link function.
@@
  • Gradient-based Training of NNs. While NNs can easily approximate a single-index model [Bac17], the statistical complexity of gradient-based learning established in prior works scales as $n\gtrsim d^{\Theta(p)}$: in the well-specified setting, [BAGJ21] proved a sample complexity of $n=\tilde{\Theta}(d^{p-1})$ for online SGD, which is later improved to $\tilde{\Theta}(d^{p/2})$ by a smoothed objective [DNGL23]; as for the misspecified setting, [BBSS22, DKL+23] showed that $n\gtrsim d^{p}$ samples suffice, and in some cases a $\tilde{\Theta}(d^{p-1})$ complexity is achievable [AAM23, OSSW24]. Consequently, at the information-theoretic limit ($n\asymp d$), existing results can only cover the learning of low information exponent targets [AAM22, BMZ23, BES+23]. It is worth noting that this exponential dependence on $p$ also appears in the CSQ lower bounds [DLS22, AAM22], which is often considered to be indicative of the performance of SGD learning with the squared loss (see Section 2.2).
  • Statistical Query Learners. If we do not restrict ourselves to correlational queries, then (1.1) can be efficiently solved near the information-theoretic limit. Specifically, [CM20] proposed an SQ algorithm that learns any single-index polynomials using $n=\tilde{O}(d)$ samples; the key ingredient is to construct nonlinear transformations to the labels that lowers the information exponent to $2$ (similar preprocessing also appeared in context of phase retrieval [MM18, BKM+19]). This is consistent with the recently established SQ lower bound [DPVLB24], which suggests a statistical complexity independent of the information exponent $p$ when $\sigma_{*}$ is polynomial.
@@ 2.2 Can Gradient Descent Go Beyond Correlational Queries?
As discussed in Section 2.1, the gap between SQ and CSQ algorithms primarily stems from the existence of label transformations that decrease the information exponent. While such transformation cannot be utilized by a CSQ learner, [DTA+24] argued that they may arise from two consecutive gradient updates using the same minibatch. For illustrative purposes, consider an example where one neuron $f_{\boldsymbol{w}}(\boldsymbol{x})=\sigma(\langle\boldsymbol{x},\boldsymbol{w}\rangle)$ is updated by two GD steps using the same training example $(\boldsymbol{x},y)$, starting from zero initialization $\boldsymbol{w}^{0}=\mathbf{0}$ (we focus on the correlational term in the loss for simplicity):

### v2

@@ Abstract
We study the problem of gradient descent learning of a single-index target function $f_{*}(\boldsymbol{x})=\textstyle\sigma_{*}\mathopen{}\mathclose{{}\left(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle}\right)$ under isotropic Gaussian data in $\mathbb{R}^{d}$, where the unknown link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ has information exponent $p$ (defined as the lowest degree in the Hermite expansion). Prior works showed that gradient-based training of neural networks can learn this target with $n\gtrsim d^{\Theta(p)}$ samples, and such complexity is predicted to be necessary by the correlational statistical query lower bound. Surprisingly, we prove that a two-layer neural network optimized by an SGD-based algorithm (on the squared loss) learns $f_{*}$ with a complexity that is not governed by the information exponent. Specifically, for arbitrary polynomial single-index models, we establish a sample and runtime complexity of $n\simeq T=\Theta(d\!\cdot\!\mathrm{polylog}d)$, where $\Theta(\cdot)$ hides a constant only depending on the degree of $\sigma_{*}$; this dimension dependence matches the information theoretic limit up to polylogarithmic factors. More generally, we show that $n\gtrsim d^{(p_{*}-1)\vee 1}$ samples are sufficient to achieve low generalization error, where $p_{*}\leq p$ is the generative exponent of the link function. Core to our analysis is the reuse of minibatch in the gradient computation, which gives rise to higher-order information beyond correlational queries.
@@ 1 Introduction
Single-index models are a classical class of functions that capture low-dimensional structure in the learning problem. To efficiently estimate such functions, the learning algorithm should extract the relevant (one-dimensional) subspace from high-dimensional observations; hence this problem setting has been extensively studied in deep learning theory [BL20, BES+22, BBSS22, MHPG+23, MZD+23, WWF24], to examine the adaptivity to low-dimensional targets and benefit of representation learning in neural networks (NNs) optimized by gradient descent (GD). In this work we study the learning of a single-index target function under isotropic Gaussian data:
@@
where $\varsigma_{i}$ is i.i.d. label noise, $\boldsymbol{\theta}\in\mathbb{R}^{d}$ is the direction of index features, and we assume the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ has information exponent $p\in\mathbb{N}_{+}$ defined as the index of the first non-zero coefficient in the Hermite expansion (see Definition 1).
Equation (1.1) requires the estimation of the one-dimensional link function $\sigma_{*}$ and the relevant direction $\boldsymbol{\theta}$; it is known that learning is information theoretically possible with $n\gtrsim d$ training examples [DH24, DPVLB24]. Indeed, when $\sigma_{*}$ is polynomial, such statistical complexity can be achieved up to logarithmic factors by a tailored algorithm that exploit the structure of low-dimensional target [CM20]. On the other hand, for gradient-based training of two-layer NNs, existing works established a sample complexity of $n\gtrsim d^{\Theta(p)}$ [BAGJ21, BBSS22, DNGL23], which presents a gap between the information theoretic limit and what is computationally achievable by (S)GD. Such a gap is also predicted by the correlational statistical query (CSQ) lower bound [DLS22, AAM23], which roughly states that for a CSQ algorithm to learn (isotropic) Gaussian single-index models using less than exponential compute, a sample size of $n\gtrsim d^{p/2}$ is necessary.
Although CSQ lower bounds are frequently cited to imply a fundamental barrier of learning via SGD (with the squared/correlation loss), strictly speaking, the CSQ model does not include empirical risk minimization with gradient descent, due to the non-adversarial noise and existence of non-correlational terms in the gradient computation. Very recently, [DTA+24] exploited higher-order terms in the gradient update arising from the reuse of the same training data, and showed that for certain link functions with high information exponent ($p>2$), two-layer NNs may still achieve weak recovery (i.e., nontrivial overlap with $\boldsymbol{\theta}$) after two GD steps with $\Theta(d)$ batch size. While this presents evidence that GD-trained NNs can learn $f_{*}$ beyond the sample complexity suggested by the CSQ lower bound, the weak recovery statement in [DTA+24] may not translate to statistical guarantees; moreover, the class of functions where SGD can achieve vanishing generalization error is not fully characterized, as only a few specific examples of link functions are discussed.
@@
Figure 1: We train a ReLU NN (3.1) with $N=1024$ neurons using SGD (squared loss) with step size $\eta=1/d$ to learn a single-index target $f_{*}(\boldsymbol{x})=\mathsf{He}_{3}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)$; heatmaps are values averaged over 10 runs. $(a)$ online SGD with batch size $B=8$; $(b)$ GD on the same batch of size $n$ for $T=2^{14}$ steps. For online SGD we only report weak recovery (i.e., averaged overlap between neuron $\boldsymbol{w}$ and target $\boldsymbol{\theta}$) since the test error does not drop.
Given the existence of (non-NN) algorithms that learn any single-index polynomials in $n=\tilde{O}(d)$ samples [CM20] regardless of the information exponent $p$, and more generally, non-CSQ algorithms with a sample complexity surpassing the CSQ lower bound [DPVLB24], it is natural to ask if gradient-based training of NNs can achieve similar statistical efficiency for this function class. Motivated by observations in [DTA+24] that SGD with reused data may break the “curse of information exponent”, we aim to address the question:
Can NN optimized by SGD with reused batch learn single-index $f_{*}$ beyond the CSQ lower bound? And for polynomial $\sigma_{*}$, can learning succeed near the information-theoretic limit $n\simeq d$?
Empirically, the separation between one-pass (online) and multi-pass SGD is clearly observed in Figure 1, where we trained the same two-layer ReLU neural network to learn a single-index polynomial with information exponent $p=3$. We see that SGD with reused data (Figure 1(b)) reaches low test error using roughly $n\simeq d$ samples, whereas online SGD fails to achieve even weak recovery with much larger sample size $n=\Omega(d^{2})$. Our main contribution is to establish this improved statistical complexity for two-layer NNs trained by a variant of SGD with reused training data.
@@ 1.1 Our Contributions
We answer the above question in the affirmative by showing that SGD training (with the squared loss) on a natural class of shallow NNs can achieve small generalization error using polynomial compute and a sample complexity that is not governed by the information exponent, if we employ a layer-wise optimization procedure (analogous to that in [BES+22, DLS22, AAM23]) and reuse of the same minibatch. The core insight is that SGD can implement a full statistical query (SQ) algorithm that goes beyond CSQ, despite the correlational structure of the squared loss. Our main finding is summarized by the following theorem.
@@
A shallow NN with $N=\tilde{\Theta}_{d}(1)$ neurons can learn arbitrary single-index models up to small population loss: $\mathbb{E}_{\boldsymbol{x}}[(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]=o_{d,\mathbb{P}}(1)$, if we employ an SGD-based algorithm (with reused training data) to minimize the squared loss objective, with a sample and runtime complexity of $n,T=\tilde{\Theta}_{d}(d^{(p_{*}-1)\vee 1})$, where $p_{*}$ is the generative exponent of the link function $\sigma_{*}$.
Note that the generative exponent [DPVLB24] is defined as the minimum information exponent of the link function $\sigma_{*}$ after arbitrary $L^{2}$ transformation, and hence by definition $p_{*}\leq p$ (equality is achieved by the identity transformation). We make the following remarks on our main result.
  • We know that $p_{*}\leq 2$ for arbitrary polynomial link functions. Therefore, the theorem suggests that NN + SGD with reused batch can learn single-index polynomials with a sample complexity $n=\tilde{O}_{d}(d)$ which is information theoretically optimal up to polylogarithmic factors, hence matching the efficiency of SQ algorithms tailored for low-dimensional polynomial regression [CM20].
  • For non-polynomial $\sigma_{*}$ with high generative exponent $p_{*}>2$, our sample complexity $n\gtrsim d^{p_{*}-1}$ can be interpreted as an SQ version of the online SGD result in [BAGJ21]. Since the information exponent $p$ can be arbitrarily larger than the generative exponent $p_{*}$, our main theorem disproves a conjecture in [AAM23] stating that $n\asymp d^{p/2}$ is the optimal sample complexity for empirical risk minimization with SGD on the squared loss / correlation loss.
  • A key observation in our analysis is that with suitable activation function, SGD with reused batch can go beyond correlational queries and implement (a subclass of) SQ algorithms. This enables polynomial transformations to the labels that reduce the information exponent, and therefore optimization can escape the high-entropy “equator” at initialization in polylogarithmic time.
Upon completion of this work, we became aware of the preprint [ADK+24] showing weak recovery (for polynomial targets with $p_{*}\leq 2$) with similar sample complexity, also by exploiting the reuse of training data. Our work was conducted independently and simultaneously.
@@
Figure 2: Complexity of learning single-index model where the link function $\sigma_{*}$ is a degree-$q$ polynomial with information exponent $p$. For the CSQ lower bound, we translate the tolerance to sample complexity using the i.i.d. concentration heuristic $\tau\approx n^{-1/2}$. We restrict ourselves to algorithms using polynomial compute; this excludes the sphere-covering procedure in [DPVLB24] or exponential-width neural network in [Bac17, TS24].
@@ 2.1 Complexity of Learning Single-index Models
We aim to learn a single-index model (1.1) where the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ has information exponent $p$ defined as follows [DH18, BAGJ21].
Definition 1 (Information exponent).
Let $\{\mathsf{He}_{j}\}_{j=0}^{\infty}$ denote the normalized Hermite polynomials. The information exponent of $g\in L^{2}(\gamma)$, denoted by $\mathrm{IE}(g):=p\in\mathbb{N}_{+}$, is the index of the first non-zero Hermite coefficient of $g$, that is, given $g(z)=\sum_{i=0}^{\infty}\alpha_{i}\mathsf{He}_{i}(z)$, $p:=\min\{i\!>\!0:\alpha_{i}\!\neq\!0\}$.
By definition, when $\sigma_{*}$ is a degree-$q$ polynomial, we always have $p\leq q$. Note that $f_{*}$ contains $\Theta(d)$ parameters to be estimated, and hence information theoretically $n\gtrsim d$ samples are both sufficient and necessary for learning [MM18, BKM+19, DPVLB24]; however, the sample complexity achieved by different (polynomial time) algorithms depends on structure of the link function.
@@
  • Gradient-based Training of NNs. While NNs can easily approximate a single-index model [Bac17], the sample complexity of gradient-based learning established in prior works typically scales as $n\gtrsim d^{\Theta(p)}$: in the well-specified setting, [BAGJ21] proved a sample complexity of $n=\tilde{\Theta}(d^{p-1})$ for online SGD, which is later improved to $\tilde{\Theta}(d^{p/2})$ by a smoothed objective [DNGL23]; as for the misspecified setting, [BBSS22, DKL+23] showed that $n\gtrsim d^{p}$ samples suffice, and in some cases a $\tilde{\Theta}(d^{p-1})$ complexity is achievable [AAM23, OSSW24a]. Consequently, at the information-theoretic limit ($n\asymp d$), existing results can only cover the learning of low information exponent targets [AAM22, BMZ23, BES+23]. This exponential dependence on $p$ also appears in the CSQ lower bounds [DLS22, AAM22], which is often considered to be indicative of the performance of SGD learning with the squared loss (see Section 2.2).
Statistical Query Learners.
If we do not restrict ourselves to correlational queries, the sample complexity of learning (1.1) can be drastically improved. Specifically, for polynomial $\sigma_{*}$, [CM20] gave an SQ algorithm that achieves low generalization error in $n=\tilde{O}(d)$ samples, which is near the information-theoretic limit; the key ingredient is to construct nonlinear transformations to the labels that lowers the information exponent to $2$; similar preprocessing also appeared in context of phase retrieval [MM18, BKM+19]. Such transformations do not belong to CSQ, but can be utilized by a full SQ learner to enhance the statistical efficiency. Recently, [DPVLB24] introduced the generative exponent which governs the complexity of SQ algorithms.
Definition 2 (Generative exponent).
The generative exponent (GE) of $g\in L^{2}(\gamma)$ is defined as the lowest information exponent (IE) after arbitrary $L^{2}$ transformation, that is,
$p_{*}=:\mathrm{GE}(g)=\inf_{\mathcal{T}\in L^{2}(P_{y})}\mathrm{IE}(\mathcal{T}\circ g).$ (2.1)
The generative exponent is the smallest information exponent obtained by all possible label transformations. By definition we always have $p^{*}\leq p$, and the gap between the two indices can be arbitrarily large; for example, for the Hermite polynomials we have $\mathrm{IE}(\mathsf{He}_{k})=k$ whereas $\mathrm{GE}(\mathsf{He}_{k})\leq 2$.
[DPVLB24] established a sample complexity lower bound of $n=\Omega(d^{\nicefrac{{p_{*}}}{{2}}\vee 1})$ for full SQ learners with polynomial compute (assuming $\tau\approx n^{-1/2}$), and obtained matching upper bound by a tensor partial-trace algorithm. Our goal is to show that SGD training of two-layer neural network can also achieve a sample and runtime complexity that scales with $n\simeq d^{\Theta(p_{*})}$, where the dimension dependence is governed by the generative exponent $p_{*}$ instead of the information exponent $p$.
@@ 2.2 Can Gradient Descent Go Beyond Correlational Queries?
As previously discussed, the gap between SQ and CSQ algorithms primarily stems from the existence of label transformations that decrease the information exponent. While such transformation cannot be utilized by a CSQ learner, [DTA+24] argued that they may arise from two consecutive gradient updates using the same minibatch. For illustrative purposes, consider one neuron $f_{\boldsymbol{w}}(\boldsymbol{x})=\sigma(\langle\boldsymbol{x},\boldsymbol{w}\rangle)$ updated by two GD steps using the same data point $(\boldsymbol{x},y)$, starting from zero initialization $\boldsymbol{w}^{0}=\mathbf{0}$ (we focus on the correlational term in the loss for simplicity):